OpenStudy (anonymous):

What is the vertex in the quadratic function y= -4x^2+8x-3? Write answer as an ordered pair.

OpenStudy (johnweldon1993):

Well you have an equation in the form \[\large ax^2 + bx + c\] To find the x-coordinate of the vertex...you use the equation \[\large \frac{-b}{2a} \] What do you get when you do that?

OpenStudy (kenljw):

y = -4x^2+8x-3 y = -4 (x^2 - 2x)-3 y = -4 ((x - 1) ^2 -1)-3 y = -4 (x - 1) ^2 +1 Vertex (1, 1)

OpenStudy (zzr0ck3r):

no reason to change form when you have -b/(2a)

OpenStudy (kenljw):

True, but you should practice you algebra instead of memorizing esp in the beginning. This form gives more than just vertex